OpenStudy (anonymous):

At a particular time, a building casts a shadow 34 ft long on the horizontal ground. At the same time, a vertical pole 5 ft high casts a shadow 9 ft long. Calculate the height of the building (in feet). (Express your answer accurate to one decimal place.)

OpenStudy (anonymous):

Utilizing triangle similarity: \[ \frac 59 = \frac x{34} \implies x=34 \times\frac 59= 18.8889 \approx 18.9 \]

OpenStudy (alexwee123):

dang

OpenStudy (anonymous):

Dang dang!

OpenStudy (alexwee123):

:(

OpenStudy (anonymous):

your solution matches my solution. thank you!